### v2

# Accelerated Rotation-Invariant Convolution for UAV Image Segmentation

Abstract
Rotation invariance is essential for precise object level segmentation in UAV aerial imagery, where targets can have arbitrary orientations and exhibit fine scale details. Conventional segmentation architectures like UNet rely on convolution operators that are not rotation invariant, leading to degraded segmentation accuracy across varying viewpoints. Rotation invariance can be achieved by expanding the filter bank across multiple orientations; however, this significantly increases computational cost and memory requirement. In this paper, we introduce a GPU optimized rotation invariant convolution framework that eliminates the traditional data lowering (im2col) step required for matrix multiplication based convolution. By exploiting structured data sharing among symmetrically rotated filters, our method achieves multi orientation convolution with greatly reduced memory requirements and computational redundancy. We further generalize the approach to accelerate convolution with arbitrary (non symmetric) rotation angles. Integrated into a UNet segmentation model, the framework yields up to a 5.7% improvement in accuracy over the non rotation aware baseline. Across extensive benchmarks, the proposed convolution achieves 20–57% faster training and 15–45% lower energy consumption than cuDNN, while maintaining accuracy comparable to state-of-the-art rotation invariant methods. Because the scatter based operator greatly reduces intermediate feature dimensionality, the efficiency of our design also enables practical sixteen orientation convolution and pooling, yielding further accuracy gains that are infeasible for conventional rotation invariant implementations. These results demonstrate that the proposed method provides an effective and highly efficient alternative to existing rotation invariant convolution frameworks.
I Introduction
Robust semantic segmentation in high‑resolution aerial images requires rotation invariance, because objects of interest—whether vegetation, buildings, vehicles, or other land‑surface features—can occur at any orientation and often contain complex fine‑scale detail. Yet the convolution operator used in mainstream deep learning segmentation architectures (e.g., U‑Net) is translation equivariant but not rotation invariant, leaving performance sensitive to object orientation. Convolution is fundamental in deep learning due to its ability to efficiently extract spatial and hierarchical features from input data [cnn-survey]. Convolution layers routinely account for the bulk of total FLOPs, memory traffic, and energy use in convolution‑based networks. They therefore dominate training time and bound inference throughput, making their optimization a central focus of prior works [FFT1], [winograd], [10286398].
These costs are further amplified when pursuing rotation‑invariant convolution, which typically requires evaluating kernels across multiple orientation samples. Despite the additional computational burden, achieving rotation invariance is essential for many aerial remote sensing applications—including satellite imagery [satellite2, satellite3, satellite4], UAV imaging [UAV1], and geospatial image localization [geo1]—because feature extraction must remain consistent under arbitrary transformations of the target object in the image, including rotations; this need is further underscored by recent UAV studies in geology (dual-spectrum hot-spring fluid segmentation) [YI2025661], urban scene segmentation (UAVformer) [YI2023109019], and UAV visual perception (CCTseg) [YI2023112612] under varying altitudes and viewpoints. The method in [mitton2021rotation] embeds rotation equivariance into U-Net for UAV image segmentation, achieving a 7% improvement in deforestation segmentation accuracy.
The extracted feature maps are rotation equivariant, i.e., feature maps whose outputs are rotated in a predictable way, as we rotate the input. A special case of equivariance is invariance [harmonic-cnn], where the output values of the feature map are not affected by rotations of the input. In essence, the rotation invariance can be achieved by directly encoding rotation equivariance into the convolution layer followed by a global pooling operation. Multiple rotated versions of the canonical filter are applied at the same input to generate rotation equivariant feature maps. For instance, Cohen & Welling [G] propose a group convolution, code-name G-convolution, in which symmetric rotations (rotation angle is multiple of 90 degrees) are applied to the same filter, generating a rotation equivariant feature map. However, such rotation equivariant convolution introduces significant overhead on the computation. For example, with G-convolution, the computational overhead increases by 4 times compared to conventional convolution in 2D space, since four symmetric rotations are applied at each filter. The computational overhead possibly increases by 24 times in 3D space because there are 24 symmetric transformations available in 3D space [CubeNet1]. The number of symmetric rotations and therefore the computational overhead increases along with increasing dimension of the input data. As reported in prior works [11125636, yang2025group], the increased number of symmetric rotations leads to substantial additional feature computations, making the training of rotation-invariant convolutional networks significantly slower than that of conventional convolutional models. In this paper, we show how the proposed convolution operation can be used to accelerate rotation invariant convolution by eliminating the data lowering step and exploiting filter weight sharing.
Modern Graphics Processing Units (GPUs) play a crucial role in deep learning, providing the computational power necessary for training deep neural networks. Their parallel processing capabilities enable efficient handling of large-scale matrix operations, significantly accelerating deep learning workloads. Matrix multiplication is an efficient method for implementing convolution on GPU, as demonstrated by [cuDNN]. This requires transforming the input image into matrices that are suitable for fast multiplication. This transformation can be achieved by lowering the input data (im2col) with duplication. The GPU implementation of the proposed method in this paper is also based on fast GPU matrix multiplication, but without data lowering.
As described in the previous work [manduhu2025airbornesensedetectdrones], conventional (gather‑style) convolution accumulates contributions at each output location: every neighboring input pixel is multiplied by its corresponding kernel weight, and the results are summed into the central output position. In contrast, our scatter convolution reverses this data flow: each input pixel is multiplied by a bank of filter weights and the resulting products are scattered directly to their destination output locations. We originally introduced this formulation to support sparse convolutions on irregularly sampled LiDAR data. In this paper, we generalize the scatter approach to (i) dense, standard grid convolution and (ii) rotation‑invariant convolution. The scatter mapping removes the im2col‑style data‑lowering step typically required by matrix multiplication based implementations, reducing memory traffic and simplifying the kernel. Moreover, because the scatter mapping exposes weight sharing across symmetric rotations, intermediate results can be reused—collapsing the number of distinct multiplications to one per symmetry group and substantially lowering computational cost.
While computational efficiency is an important consideration, achieving rotation invariance in UAV image segmentation inherently involves a trade-off between segmentation accuracy and computational efficiency. Increasing the number of modeled orientations generally improves accuracy by enhancing rotational robustness and feature expressiveness; however, this improvement comes at the cost of increased computation, memory traffic, and intermediate feature-map growth. As demonstrated in our experiments, moderate orientation resolutions (e.g., eight orientations) provide a favorable balance by significantly improving segmentation accuracy while remaining computationally feasible. In contrast, further increasing the orientation resolution leads to prohibitively large intermediate feature maps for conventional rotation-invariant implementations, making them infeasible to execute within GPU memory constraints. This work therefore focuses not only on improving computational efficiency, but also on enabling practical high-resolution rotation handling by reducing redundant computation and memory overhead, allowing finer orientation sampling under realistic GPU constraints.
Our main contributions are as follows:
  • GPU-optimized single-orientation convolution: Since a single-orientation convolution reduces to the standard convolution, we propose an efficient GPU implementation of standard convolution using matrix multiplication without data lowering (im2col), which significantly reduces computational operations and memory access overhead.
  • Symmetric rotation optimization: For symmetric rotations (90°, 180°, 270°, and 360°), we exploit data sharing via scatter operations, allowing each multiplication to be computed once and reused across all symmetric orientations.
  • Arbitrary rotation acceleration: The framework is further generalized to support arbitrary rotation angles, achieving efficient convolution under continuous rotational transformations.
  • Comprehensive evaluation: Integrated into a UNet model, the proposed convolution yields up to a 5.7% improvement over the non rotation aware baseline, while achieving 20–57% faster training and 15–45% lower energy consumption than cuDNN across all evaluated workloads. Its efficient scatter based design also enables practical sixteen orientation convolution, providing additional gains unattainable with conventional rotation invariant methods.
II Related Work
II-A Segmentation of remote sensing imagery
Early semantic segmentation models were largely based on convolutional neural networks (CNNs) [7789580]. A key milestone was the fully convolutional network (FCN) introduced in [FCN], which adopts an encoder–decoder architecture composed entirely of convolutional layers, without any fully connected layers. Building on this foundation, [rs13234902] proposed an improved FCN framework for remote sensing image segmentation, incorporating multi-scale feature extraction, enhanced encoder–decoder connections, and refined upsampling to better handle complex, high-resolution imagery. To obtain larger receptive fields without lowering image resolution via downsampling, DeepLab was introduced in [deeplabv1], utilizing atrous convolutions to achieve dense, multi-scale feature extraction. The work in [10608051] further improves DeepLab by aggregating multi-scale features to enhance aerial image segmentation. Many early segmentation models did not fully exploit low-level features in the decoder, resulting in the loss of fine spatial detail. U-Net [ronneberger2015u] addresses this limitation by introducing skip connections that link encoder and decoder features, enabling more effective fusion of detailed and high-level information. U-Net and its variants have also been widely adopted for UAV and aerial image segmentation, as in UVid-Net for UAV video segmentation [9392319] and Res-U-Net-based coastal boundary detection from high-resolution UAV imagery [wang2025multi].
Attention-based architectures have also been used for aerial image segmentation. SegHSI [10751785] adopts a CNN-based backbone augmented with cluster attention to model spectral–spatial correlations in hyperspectral imagery. The work in [10422983] employs a Transformer–CNN hybrid encoder–decoder for UAV forestry segmentation, enabling accurate delineation of diseased pine canopies in high resolution imagery.
II-B Rotation invariance in remote sensing imagery
Rotation invariance (or equivariance) is crucial for accurate object-level segmentation in UAV imagery, as targets often appear in arbitrary orientations. The work in [isprs-annals-V-1-2022-15-2022] introduces a CNN-based framework for vehicle instance segmentation in UAV images that uses rotated bounding boxes instead of standard axis-aligned boxes to better match the true orientation of vehicles. The method in [9466361] learns rotation-invariant deep embeddings by aligning features from multiple rotated views of the same image, enabling robust analysis of remote sensing images under arbitrary orientations. FRINet, introduced in [10339376], is built on CNN feature extractors and incorporates a rotation-adaptive matching mechanism to enable few-shot, rotation-invariant segmentation of aerial images. The paper [mo2024achieving] proposes a mechanism-assured convolution that embeds rotation invariance directly into the operation itself, avoiding reliance on data augmentation and ensuring consistent feature responses under rotated inputs. RIC-CNN [mo2024ric] achieves rotation invariance by embedding polar coordinate information into convolution operations, enabling the network to generate stable feature representations under arbitrary rotations. The method in [mitton2021rotation] embeds rotation equivariance into UNet by using the discrete rotation group $C_{8}$ ($45^{\circ}$ increments) as the symmetry group, enabling stable feature representations that improve both deforestation segmentation and driver classification under arbitrary image orientations. In this paper, we introduce a computationally efficient rotation-invariant convolution and incorporate it into a U-Net architecture for semantic segmentation of UAV imagery.
II-C Single-orientation rotation convolution
As single-orientation convolution reduces to standard convolution, we review efficient implementations of standard convolution in this section. Several FFT-based approaches [FFT1, FFT2, FFT3, FFT4] have been proposed to reduce the computational cost for standard convolution on different hardware. FFT-based approaches can significantly lower the work complexity of convolution, but it is only effective for convolution with larger filters [cuDNN].
Lavin & Gray [Lavin] proposed a method based on the Winograd algorithm [winograd] for convolution with small filters. The key idea of Winograd-based convolution is similar to the one based on FFT which performs multiplications in the complex domain. However, unlike FFT-based convolution, the Winograd-based convolution operates on real numbers, thus requiring fewer operations. For a filter with larger size, Winograd-based convolution shows poor performance, since it needs to perform a significant amount of extra arithmetic such as additions and data transformations.
To reduce the substantial memory usage in convolution with matrix multiplication, Lu et al. [lu2023im2win] proposed new data-lowering techniques. However, their method still requires data duplication. In [10.5555/3213069.3213072], Chang et al. proposed an approach to improve memory efficiency in convolution operations by effectively utilizing the on-chip memory of GPUs. In [9235051], a matrix multiplication-based approach without data lowering is proposed. However, it requires recomputing writing positions, preventing direct pointwise operations.
II-D Rotation-equivariant convolution for multi-orientation rotations
A general theory of an equivariant CNN can be found in [cohen2018general]. The mathematical formulation of equivariance through a weight sharing scheme is exhibited in [equi-theory].
G-convolution [G] applies a group of filters to each feature map, where each filter is a transformed version of the original filter. Two different groups are proposed in [G], code-name $p4$ group and $p4m$ group. The $p4$ group consists of 4 filters where each filter is obtained by rotating the original filter by 0 degrees, 90 degrees, 180 degrees and 270 degrees, respectively. The $p4m$ group consists of 8 filters which are obtained by combining $p4$ group and reflection transformations. The convolution with $p4$ group or $p4m$ group will increase the computational overhead by 4 or 8 times. There are multiple works [CubeNet1, CubeNet2] applying the G-convolution on 3D volume data. In [CubeNet1], different groups code-name Cube Group, Klein’s Four-group and Tetrahedral Group are proposed. Among these groups, the largest group is the Cube Group which consists of 24 filters resulting in an increase of computational overhead by 24 times. Obviously, the number of filters in a group will increase along with increasing dimension of the input data and therefore the computational overhead. Hexagonal group convolution [hexa] rotates a filter by any multiple of 60 degrees, without interpolation. They define the following two hexagonal group convolutions: groups $p6$ and $p6m$, which contain integer translations (orientation cycling), rotations with multiples of 60 degrees, and mirroring for $p6m$. The paper [G-cnn10] proposes Gauge-equivariant convolutional network for manifolds, which uses Hexagonal group convolution to implement the weight sharing scheme for Gauge-equivariant. The work in [shallowNet] proposes a shallow network which performs convolution with rotated versions of each canonical filter, where the rotation is with arbitrary angle. The convolution is followed by a global pooling over orientations. The number of rotations applied is 32 i.e. the angle increment between successive two rotations is $\frac{\pi}{16}$. Bicubic interpolation is used for sampling after rotation.
Oriented Response Network (ORN) [orn] introduces active rotating filters that generate orientation-aware feature maps, allowing CNNs to model rotations explicitly and achieve rotation invariance or equivariance without heavy data augmentation. The work in [e2cnn] provides a unified framework for designing steerable CNNs that are equivariant to all $E(2)$ transformations and offers the widely used E2CNN library for generating rotation-equivariant features.
Filters with steerability can be reconstructed at any orientation as a finite, linear combination of base filters. This eliminates the need to learn multiple filters at different orientations. Harmonic Network proposed in [harmonic-cnn] achieves rotation equivariance in continuous rotations by learning coefficients of base filters, where these base filters are used to form the filters (steerable) of the CNN model. In the network proposed in [G-cnn7], rotation equivariance is guaranteed by using the G-convolution in which the number of rotation is more than four. The rotated filter (steerable) in a group is constructed by manipulating coefficients of the base filter. In this way, it avoids the use of interpolation which usually introduces artifacts in the rotated filter. The network shown in [gabor] achieves rotation equivariance by multiplying each convolutional filter by several oriented Gabor filters [weldon1996efficient]. In the back propagation stage, only the convolutional filters are updated. The method in [cheng2018rotdcf] reduces the number of parameters and computational complexity of rotation-equivariant CNNs by decomposing convolutional filters (steerable) under joint steerable bases over space and rotations simultaneously. Meanwhile, the performance is preserved.
Rotation equivariance typically leads to a rapid increase in feature dimensionality as the number of sampled orientations grows. To tackle this problem of exploding dimensionality, a rotation equivariant convolution is usually followed by an orientation pooling operation, after which the layer becomes rotation invariant. The orientation pooling can be implemented in two different ways: point-wise average pooling [featureTrans] or point-wise max pooling [shallowNet].
In this paper, we describe how our proposed convolution operation is applied to accelerate a rotation invariant convolution layer, which combines symmetric rotation and steerable transformations.
III Methodology
UAV aerial image segmentation presents unique challenges compared with ground-level imagery, as target objects often appear at arbitrary orientations due to the aerial imaging perspective. Conventional convolution operators are sensitive to such orientation variations, which can lead to degraded segmentation accuracy. The methodology presented in this section introduces a series of convolution formulations that are specifically designed as drop-in replacements for standard convolution layers in semantic segmentation networks.These operators form the computational backbone of the proposed UAV segmentation framework, enabling rotation-aware feature extraction while maintaining efficiency suitable for large-scale aerial imagery. Their practical effectiveness is demonstrated through integration into a U-Net architecture and validated in the experimental section.
III-A Single orientation convolution formulation
In the single-orientation case, the operation reduces to standard convolution because only one fixed kernel orientation is applied and no orientation-dependent processing is required. In standard convolution, each filter element is multiplied by its overlapping input element, and the resulting products are summed to produce the output value at the corresponding spatial location, as illustrated in Fig. 1(a). Alternatively, as demonstrated in our previous work [manduhu2025airbornesensedetectdrones], convolution can also be performed using a scatter-based approach (which we employed to implement sparse convolution). In this approach, each input pixel is multiplied by different kernel elements, and the resulting values are added to their corresponding neighboring output positions, as shown in Fig. 1(b). This distributes computation across output locations rather than accumulating it at a single pixel, enabling convolution without data lowering while benefiting from the high-throughput matrix multiplication capabilities of modern GPUs. In this paper, we extend the scatter-based approach to support dense convolution.
Figure 1: (a) Standard convolution where $x_{i},w_{i},y_{i}$ represent input, filter and output, respectively. (b) Convolution with scatter operation where each input element is multiplied by each filter element and the results are added with those of neighboring elements.
Let the input be $\mathbf{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$ and the filters $\mathbf{W}\in\mathbb{R}^{C_{\mathrm{out}}\times C_{\mathrm{in}}\times K_{h}\times K_{w}}$, where $C_{\mathrm{in}}$ and $C_{\mathrm{out}}$ denote the numbers of input channels and output channels, respectively, and $H$ and $W$ represent the height and width of the input feature map, while $K_{h}$ and $K_{w}$ denote the kernel height and width. For valid convolution, the output spatial size is
$H^{\prime}=H-K_{h}+1,\qquad W^{\prime}=W-K_{w}+1.$
The convolution output is
$\begin{split}Y_{c_{o},h,w}=\sum_{c_{i}=0}^{C_{\mathrm{in}}-1}\sum_{i=0}^{K_{h}-1}\sum_{j=0}^{K_{w}-1}W_{c_{o},c_{i},i,j}\;X_{c_{i},\,h+i,\,w+j},\\ h\in\{0,...,H^{\prime}-1\}\\ w\in\{0,...,W^{\prime}-1\}\\ c_{o}\in\{0,...,C_{out}-1\}\end{split}$ (1)
which is the standard sliding-window inner product formulation. To convert convolution into matrix multiplication, every $K_{h}\times K_{w}$ receptive field from $\mathbf{X}$ is flattened into a column, forming $\mathbf{X}_{\mathrm{col}}\in\mathbb{R}^{(C_{\mathrm{in}}K_{h}K_{w})\,\times\,(H^{\prime}W^{\prime})}.$ Formally,
$X_{\mathrm{col}}(k,t)=X(c_{i},\,h+i,\,w+j),$ (2)
where
$k=c_{i}K_{h}K_{w}+iK_{w}+j,\qquad t=hW^{\prime}+w,$
and $h$ and $w$ enumerate output spatial locations ($h\in\{0,\ldots,H^{\prime}-1\},\;w\in\{0,\ldots,W^{\prime}-1\}$), while $i$ and $j$ enumerate positions inside the $K_{h}\times K_{w}$ receptive field ($i\in\{0,\ldots,K_{h}-1\},\;j\in\{0,\ldots,K_{w}-1\}$). Thus $X(c_{i},\,h+i,\,w+j)$ selects the input element covered by the kernel at offset $(i,j)$ for input channel $c_{i}$. Each column corresponds to one sliding patch, and each row indexes one element in the flattened receptive field.
Each filter kernel is flattened into a row vector:
$\mathbf{W}_{\mathrm{row}}\in\mathbb{R}^{C_{\mathrm{out}}\times(C_{\mathrm{in}}K_{h}K_{w})},$
with
$W_{\mathrm{row}}(c_{o},k)=W(c_{o},c_{i},i,j).$
The convolution can now be computed as a single matrix multiplication:
$\mathbf{Y}_{\mathrm{col}}=\mathbf{W}_{\mathrm{row}}\,\mathbf{X}_{\mathrm{col}}.$
Clearly, flattening the receptive fields of $\mathbf{X}$ causes each input element to be replicated $K_{h}K_{w}$ times, i.e., $\mathbf{X}_{\mathrm{col}}$ contains up to $K_{h}K_{w}$ copies of each element of $\mathbf{X}$. Consequently, the number of multiplications increases by a factor of $K_{h}K_{w}$.
For a single-channel input and a single filter, the general formula of the convolution with scattering operation is given as follows:
[…]
For $C_{\mathrm{in}}$ input channels and $C_{\mathrm{out}}$ output channels, the convolution with scattering can be expressed as follows:
$\begin{split}Y_{c_{o},i-m+\lfloor K_{h}/2\rfloor,j-n+\lfloor K_{w}/2\rfloor}\;+=\\ \sum_{c_{i}=0}^{C_{in}-1}{X_{c_{i},h,w}\times W_{c_{o},c_{i},m,n}}\\ m\in\{0,...,K_{h}-1\},n\in\{0,...,K_{w}-1\}\\ h\in\{0,...,H^{\prime}-1\},w\in\{0,...,W^{\prime}-1\}\\ c_{o}\in\{0,...,C_{out}-1\}\end{split}$ (4)
As shown on the right-hand side of Formula (4), the matrix multiplication is used to perform channel-wise multiplication and summation. The results are then added to the corresponding output positions following the scatter rules. This approach eliminates the need for data duplication when performing convolution using matrix multiplication.
The pseudocode for single-channel scatter convolution is shown in Algorithm 1. The number of multiplications and additions required is the same as standard convolution, it is $O(H\times W\times K_{h}\times K_{w})$.
Input: An image $X(i,j)$ of size $H\times W$ and a filter $W(m,n)$ of size $K_{h}\times K_{w}$
Output: A new image $Y(x,y)$ after convolution
for $i\leftarrow 0$ to H1H-1 do
 for $j\leftarrow 0$ to W1W-1 do
    for $m\leftarrow 0$ to $K_{h}-1$ do
       for $n\leftarrow 0$ to Kw1K-w-1 do
          $x\leftarrow i-m+\lfloor K_{h}/2\rfloor$;
          $y\leftarrow j-n+\lfloor K_{w}/2\rfloor$;
          $Y(x,y)\leftarrow Y(x,y)+X(i,j)\times W(m,n)$;
         end for
      end for
   end for
end for
return $Y(x,y)$;
Algorithm 1 Convolution with scatter operation
III-B Symmetric rotation-invariant convolution formulation
Symmetric rotation equivariant can be achieved by rotating a filter under the discrete rotational symmetries of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. This set of four rotations defines the $p4$ group. A single base kernel is rotated by these four angles, producing four orientation-specific filters that capture features consistently across different orientations.
Let the four discrete rotations be indexed by $r\in\{0,1,2,3\}$, corresponding to angles $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$. Let $R_{r}$ denote the action that rotates spatial coordinates by the angle associated with $r$. For an input $\mathbf{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$ and kernels $\mathbf{W}\in\mathbb{R}^{C_{\mathrm{out}}\times C_{\mathrm{in}}\times K_{h}\times K_{w}}$, the $r$-rotated kernels are defined by
$\mathbf{W}^{(r)}_{c_{o},c_{i}}\;=\;R_{r}\!\big(\mathbf{W}_{c_{o},c_{i}}\big),\qquad\begin{subarray}{c}\displaystyle c_{o}\in\{0,\dots,C_{\mathrm{out}}-1\},\\[2.0pt] \displaystyle c_{i}\in\{0,\dots,C_{\mathrm{in}}-1\}.\end{subarray}$ (5)
For valid convolution, the output spatial size is
$H^{\prime}=H-K_{h}+1,\qquad W^{\prime}=W-K_{w}+1.$ (6)
The output has an orientation channel for each $r$:
$\begin{split}Y_{c_{o},r,h,w}&=\sum_{c_{i}=0}^{C_{\mathrm{in}}-1}\sum_{i=0}^{K_{h}-1}\sum_{j=0}^{K_{w}-1}W^{(r)}_{c_{o},c_{i},i,j}\;X_{c_{i},\,h+i,\,w+j},\\[6.0pt] &\qquad\begin{subarray}{c}\displaystyle c_{o}\in\{0,\dots,C_{\mathrm{out}}-1\},\\ \displaystyle r\in\{0,1,2,3\},\\ \displaystyle h\in\{0,\dots,H^{\prime}-1\},\\ \displaystyle w\in\{0,\dots,W^{\prime}-1\}\end{subarray}\end{split}$ (7)
While this improves rotational equivariance, it also increases computational cost because each rotated filter must be evaluated separately. To mitigate this cost, we demonstrate how $p4$ group convolution on a 2D input can be accelerated using the scatter operation. In the $p4$ group, the original filter is rotated four times at $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, respectively. Figure 2 illustrates the scatter-based convolution process for each rotated filter when applied to the input pixel $x_{13}$. It is clear the same product between a filter weight and an input element appears in all 4 convolutions albeit in different locations. For example, the multiplication $x_{13}\times w_{9}$ appears in 4 different corners of all convolutions. We can compute the multiplication $x_{13}\times w_{9}$ for the convolution with the first filter and use the result of this multiplication in the other convolutions directly. The multiplication is performed for only one filter. The index of the output location in which we add the result of the multiplication can be pre-computed for each filter. The same idea also can be applied to $p4m$ group convolution in which the original filter transformed 8 times (rotation and reflection). The multiplication is performed for the first filter and the result is reused by the other 7 filters directly.
Figure 2: Implementation of $p4$ group convolution with scatter operation where the original filter is rotated (a) $0^{\circ}$ (b) $90^{\circ}$ (c) $180^{\circ}$ (d) $270^{\circ}$.
Formally, for the four symmetric rotations $r\in\{0,1,2,3\}$, corresponding to $0^{\circ},90^{\circ},180^{\circ},$ and $270^{\circ}$, let $(m^{\prime},n^{\prime})=R_{r}(m,n)$ denote the rotated kernel coordinates obtained by applying the spatial rotation operator $R_{r}$ to the original offsets $(m,n)$. Under this rotation, the scatter-based convolution becomes
$\begin{split}Y_{c_{o},\;i-m^{\prime}+\lfloor K_{h}/2\rfloor,\;j-n^{\prime}+\lfloor K_{w}/2\rfloor}\;+=\;\\ \sum_{c_{i}=0}^{C_{\mathrm{in}}-1}X_{c_{i},h,w}\;W_{c_{o},c_{i},m,n},\\[6.0pt] \begin{subarray}{c}\displaystyle m\in\{0,\dots,K_{h}-1\},\qquad\displaystyle n\in\{0,\dots,K_{w}-1\},\\[4.0pt] \displaystyle h\in\{0,\dots,H^{\prime}-1\},\qquad\displaystyle w\in\{0,\dots,W^{\prime}-1\},\\[4.0pt] \displaystyle c_{o}\in\{0,\dots,C_{\mathrm{out}}-1\},\qquad\displaystyle r\in\{0,1,2,3\}.\end{subarray}\end{split}$ (8)
In this expression, the channel-wise multiplication and accumulation term $\sum\limits_{c_{i}=0}^{C_{\mathrm{in}}-1}X_{c_{i},h,w}\,W_{c_{o},c_{i},m,n}$ is computed only once. For each rotation $r$, this accumulated product is added to a different output location because the kernel coordinate $(m,n)$ is transformed to its rotated counterpart $(m^{\prime},n^{\prime})=R_{r}(m,n)$. Thus, only the scatter location changes across rotations, while the channel-wise multiplication and accumulation is reused for all four orientations.
A rotation-invariant response can be obtained by group pooling over the four orientations. For example, average pooling yields
$\widetilde{Y}_{c_{o},h,w}=\frac{1}{4}\sum_{r=0}^{3}Y_{c_{o},r,h,w}.$ (9)
Alternatively, max pooling over orientations may be used:
$\widetilde{Y}_{c_{o},h,w}=\max_{r\in\{0,1,2,3\}}\;Y_{c_{o},r,h,w}.$ (10)
For symmetric rotation equivariant convolution, the number of filters will increase with the increase of the dimension of the input data. Consequently, the computational overhead also increases. However, with the scatter operation, the multiplication just needs to be performed for one filter and the result of the multiplication can be reused by the other filters directly. It means the number of multiplications is a constant (with respect to the number of directions) and the convolution is only dominated by the addition operation.
Now we introduce the gradient computation for the rotation-invariant convolution. During backpropagation, the loss gradient propagates from the output toward the input. The incoming (upstream) gradient from the next layer is denoted by
$G_{c_{o},h,w}=\frac{\partial\mathcal{L}}{\partial\widetilde{Y}_{c_{o},h,w}}.$
where $\widetilde{\mathbf{Y}}$ is the rotation-invariant output feature map obtained after orientation pooling. We derive the gradients with respect to the rotation-equivariant feature maps $\mathbf{Y}_{c_{o},r,h,w}$, the input feature maps $\mathbf{X}$, and the base convolution kernel $\mathbf{W}$.
For average pooling, the gradient is distributed equally:
$\frac{\partial\mathcal{L}}{\partial Y_{c_{o},r,h,w}}=\frac{1}{4}\,G_{c_{o},h,w}.$ (11)
For max pooling, we define
$r^{\star}(c_{o},h,w)\in\arg\max_{r}Y_{c_{o},r,h,w}.$
Then the gradient passes only through the maximal rotation:
$\frac{\partial\mathcal{L}}{\partial Y_{c_{o},r,h,w}}=G_{c_{o},h,w}\;\mathbf{1}\!\left[r=r^{\star}(c_{o},h,w)\right].$ (12)
Let
$\Delta_{c_{o},r,h,w}=\frac{\partial\mathcal{L}}{\partial Y_{c_{o},r,h,w}}$
for compactness in the following derivations.
For each rotation $r\in\{0,1,2,3\}$, the gradients with respect to the input and the $r$-rotated kernel $\mathbf{W}^{(r)}$ are given by the standard convolution backward formulas:
$\frac{\partial\mathcal{L}}{\partial X_{c_{i},h,w}}^{(r)}=\sum_{c_{o}=0}^{C_{\mathrm{out}}-1}\sum_{i=0}^{K_{h}-1}\sum_{j=0}^{K_{w}-1}W^{(r)}_{c_{o},c_{i},i,j}\;\Delta_{c_{o},r,\,h-i,\,w-j},$ (13)
$\frac{\partial\mathcal{L}}{\partial W^{(r)}_{c_{o},c_{i},i,j}}=\sum_{h=0}^{H^{\prime}-1}\sum_{w=0}^{W^{\prime}-1}X_{c_{i},\,h+i,\,w+j}\;\Delta_{c_{o},r,h,w}.$ (14)
The total gradient with respect to the input feature map is the sum over the four rotated branches:
$\frac{\partial\mathcal{L}}{\partial X_{c_{i},h,w}}=\sum_{r=0}^{3}\frac{\partial\mathcal{L}}{\partial X_{c_{i},h,w}}^{(r)}.$ (15)
To obtain the gradient with respect to the unrotated base kernel $\mathbf{W}$, each rotated kernel gradient is first inverse-rotated and then accumulated:
$\frac{\partial\mathcal{L}}{\partial W_{c_{o},c_{i},i,j}}=\sum_{r=0}^{3}R_{-r}\!\left(\frac{\partial\mathcal{L}}{\partial W^{(r)}_{c_{o},c_{i},i,j}}\right)$. (16)
Here $R_{-r}$ denotes a rotation of the gradient by $r\times 90^{\circ}$, mapping each rotated kernel’s gradient back to the base orientation. The proposed scatter approach can also accelerate backpropagation, since the gradient computations themselves are convolution operations, as shown in these formulations.
III-C Parallelization of the scatter operation
The parallelization of the convolution with scatter operation is shown in Fig. 3. In each parallel step, every processor $P_{i}$ reads the corresponding input data $x_{i}$ and multiplies it with the same filter weight, then the results of those multiplications will be added to the different output locations. Multiplication with the same filter weight can avoid Write after Write hazard, if there is a synchronization between two parallel steps. For G-convolution, the multiplication is performed only once using this approach.
Figure 3: Parallel scatter operation, where each processor $P_{i}$ corresponds to each input data.
The scatter operation is a cache friendly operation and can be easily vectorized. As shown in Fig. 4, if all reading positions in the input data are in the same row and are multiplied by the same filter weight, then all writing positions in the output are always in the same row, no matter which filter weight is multiplied. Even if the filter is rotated symmetrically, all writing positions in the output are also in the same row, see Fig. 5. It meets the data locality requirement of modern SIMD architecture. On the other hand, this figure also shows us a way to reduce the number of synchronizations required. All processors multiply its own input data with the same filter weight and add the result in 4 different outputs. Obviously, before completing the last addition, all processors have no need to be synchronized since they are writing different memory positions of different outputs. We just insert a synchronization after the last addition. The number of synchronizations can be reduced 4 times.
Figure 4: Illustration of scatter operation by multiplying each input data with another filter weight. No matter which filter weight it is multiplied by, all writing positions in the output are always in the same row.
Figure 5: Output positions for the rotated filter with data reuse. No matter how the filter is rotated, all writing positions in the output are always in the same row if the processors process input data in one row.
III-D Arbitrary rotation convolution formulation
An interpolation is required for sampling after rotation with arbitrary angle. However, it can introduce artifacts that degrade the performance of rotation-invariant models. In this subsection, we propose an approach which takes the advantage of symmetric rotation and steerable filter to implement rotation-invariant convolution with filter arbitrarily rotated. The proposed method accelerates the training process while eliminating the need for interpolation.
Figure 6: Symmetric correspondents of $45^{\circ}$ rotated image in other quadrants of coordinate plane
[…]
As steerable filters can generate rotated responses without relying on interpolation, we construct the filter in the first quadrant of the coordinate plane using the following steerable filter formulation as in [93808]:
$\psi_{\theta}(x,y)=\sin(\theta)\cdot f_{x}(x,y)+\cos(\theta)\cdot f_{y}(x,y)$ (17)
Equation (17) defines the steerable filter $\psi_{\theta}(x,y)$ at a given rotation angle $\theta$, where $(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates. The functions $f_{x}(x,y)$ and $f_{y}(x,y)$ represent the learned base filters oriented along the horizontal (x-axis) and vertical (y-axis) directions, respectively. The steerable filter $\psi_{\theta}(x,y)$ is constructed as a linear combination of these directional basis filters using fixed orientation-dependent coefficients $\sin(\theta)$ and $\cos(\theta)$, enabling efficient and continuous steering of the filter to any desired angle $\theta$.
To encourage the learned base filters to exhibit steerable behavior, we add two regularization losses to the standard cross-entropy objective. The overall loss is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda_{\text{mag}}\,\mathcal{L}_{\text{mag}}+\lambda_{\text{orth}}\,\mathcal{L}_{\text{orth}},$ (18)
where $\mathcal{L}_{\text{CE}}$ denotes the task-specific cross-entropy loss, and $\lambda_{\text{mag}}$ and $\lambda_{\text{orth}}$ are scalar weighting coefficients controlling the strength of the magnitude and orthogonality regularizations, respectively.
The magnitude-matching loss encourages the two base filters to maintain similar energy levels:
$\mathcal{L}_{\text{mag}}=\frac{1}{B}\sum_{b=1}^{B}\left(\left\|\mathbf{w}^{(x)}_{b}\right\|_{2}-\left\|\mathbf{w}^{(y)}_{b}\right\|_{2}\right)^{2},$ (19)
where $\mathbf{w}^{(x)}_{b}$ and $\mathbf{w}^{(y)}_{b}$ denote the flattened kernel weights associated with the $b$-th filter in the base filters $f_{x}$ and $f_{y}$, respectively, and $B$ is the total number of filters.
The orthogonality loss penalizes the correlation between the two directional bases, promoting rotational independence:
$\mathcal{L}_{\text{orth}}=\frac{1}{B}\sum_{b=1}^{B}\left(\frac{\left\langle\mathbf{w}^{(x)}_{b},\mathbf{w}^{(y)}_{b}\right\rangle}{\left\|\mathbf{w}^{(x)}_{b}\right\|_{2}\left\|\mathbf{w}^{(y)}_{b}\right\|_{2}+\varepsilon}\right)^{2},$ (20)
where $\varepsilon\!>\!0$ ensures numerical stability during training. Together, these regularization terms guide the base filters $f_{x}$ and $f_{y}$ toward forming an approximately steerable pair that preserves orientation consistency across rotations.
IV GPU implementation
IV-A GPU implementation of single-orientation convolution with scatter operation
The convolution operation can be efficiently implemented on GPU using matrix multiplication, as shown in [cudnn-arxiv]. This approach requires transforming the input image and the convolution kernel into matrices that can exploit the fast multiplication capabilities of the GPU. This transformation involves duplicating the input data to lower its dimensionality.
In our GPU implementation of convolution with scatter operation, we use matrix multiplication to perform the channel-wise multiplication and summation of the input and a filter kernel. To achieve a higher performance, NVIDIA cuBLAS library [cublas], a library of highly optimized matrix multiplication routines, is used to perform the channel-wise multiplication and summation. The results are then scattered to the corresponding output positions according to the scattering rule. Unlike the method in [cudnn-arxiv], we do not apply any data lowering in our GPU implementation. Our GPU implementation follows gather-GEMM-scatter dataflow [torchsparse] but gather without data lowering.
Considering that cuBLAS accesses matrices in column-major order, we arrange the input image in CNHW layout and the filter in NHWC layout, as shown in Fig. 7. The resulting matrix ensures coalesced access for the parallel implementation of the scatter operation, as described in Section III-C.
Figure 7: Memory layout of the input image and filter kernel for matrix multiplication, where $P_{i,j}$ represents an image pixel, $F_{m,n}$ represents an element of the filter kernel, $c_{i,j,L}$ and $c^{\prime}_{m,n,L}$ are channel values. Only a single image and a single filter are depicted.
After performing matrix multiplication, the results need to be added to the output image using the parallel scattering operation. However, directly writing to off-chip memory (i.e., the global memory of CUDA) is time-consuming, as each element of the output image needs to be accessed $k^{2}$ times. To avoid this, we tile the output image as in Fig. 8 so that each tile can be accommodated in on-chip memory (i.e., the shared memory of CUDA). The scatter operation will be performed in on-chip memory, and the results will then be written to off-chip memory. There is an overlapping region between adjacent tiles, which ensures that elements on the borders can be scattered to neighboring positions completely.
Figure 8: Image tiling with on-chip memory of GPU, where the size of overlap between tiles is $\lfloor k/2\rfloor$.
IV-B GPU implementation of symmetric rotation invariant convolution with scatter operation
The main idea behind accelerating symmetric rotation invariant convolution using the scatter operation is the reuse of intermediate multiplication results, as described in Section III-B. The matrix-multiplication outputs are shared across all rotations, eliminating redundant computation. To make the layer rotation invariant, an orientation-pooling operation is applied afterward, which can be implemented in two ways: point-wise average pooling or point-wise max pooling.
For point-wise average pooling, each element in the result of the matrix multiplication is divided by $r$, the number of rotations, before being added to the output image. This means that the required on-chip memory size is equal to the tile size. However, for point-wise max pooling, the on-chip memory must be twice the tile size—one part for computing the current tile and another for storing the previous maximum values. After finishing additions for the current tile, the point-wise max operation is performed between the current tile and the previously stored max tile. This is necessary because max pooling is a non-linear operation, meaning the maximum value cannot be determined before the additions are completed. Additionally, the overlapping region between adjacent tiles must be extended to $\lfloor k/2\rfloor+1$ to ensure correct max pooling calculations. Only the central part of each tile, excluding $\lfloor k/2\rfloor+1$ border lines, is used in the final output. This ensures that the max values are computed correctly without missing contributions from neighboring tiles.
V Performance Evaluation
We implemented the rotation-invariant convolution using four different approaches: (1) our scatter based kernel, (2) a cuDNN based pipeline, (3) an Oriented Response Networks (ORN) implementation, and (4) an equivariant convolution implementation using the e2cnn library.
In our scatter based approach, the four symmetric rotations are processed within a single fused convolution, so each input element is multiplied only once. In contrast, the cuDNN based pipeline must treat every rotated filter as an independent kernel, performing four separate convolution and therefore four distinct multiply–accumulate passes before the results can be combined. The ORN and e2cnn implementations follow their respective rotation-equivariant formulations and are included as additional baselines for comparison. Among these, we implemented the backward pass for the scatter-based and cuDNN-based versions, as automatic gradient computation is not available for these custom rotation-invariant operations, ensuring that gradients propagate correctly through every rotated filter.
V-A Model for Evaluation
To evaluate the performance of the scatter based rotation invariant convolution in a real segmentation application, we replaced the first convolution of base convolution block in UNet [ronneberger2015u] with the rotation invariant convolution described above (see Fig. 9 for illustration). To balance performance and dimension explosion, we apply max pooling across every four symmetric rotations; for example, with eight rotations, this produces two orientation features, each corresponding to a group of four symmetric rotations. The second convolution within the base block uses a standard convolution, we found through experiments that this design improves performance.
Figure 9: Unet architecture with rotation invariant convolution block.
V-B Datasets for Training and Testing
The modified UNet were trained and tested on two datasets: an in-farm plant-segmentation dataset captured with a DJI M300 drone [dji_m300] equipped with a P1 photogrammetry camera [dji_p1], and the public Semantic Segmentation Drone Dataset [semantic_drone_dataset]. All experiments were conducted on an Ubuntu 22.04 system equipped with eight NVIDIA RTX 3090 GPUs.
V-B1 Plant segmentation dataset
Accurately measuring species composition in pasture and forage systems is fundamental for biodiversity monitoring and precision agriculture because it informs management decisions on seeding, fertilization, and grazing intensity. However, mixed‑species pastures are highly heterogeneous: the realized stand rarely matches the seed plan and can vary widely across a field and over time. This variability places stringent demands on automated, high‑resolution, plant‑level segmentation of aerial/UAV imagery. Because leaves and other canopy structures appear at arbitrary orientations in overhead views, incorporating rotation invariance into the segmentation pipeline is critical for robust species discrimination in multi‑species pastures.
We mapped pasture species composition from high‑resolution UAV imagery acquired with a DJI drone equipped with a Zenmuse P1 camera. Flights were conducted at approximately 12 m above ground level, yielding a nominal ground sampling distance (GSD) of 0.15 cm per pixel. Forward image overlap was 80%, and each P1 frame covered roughly 12.31 x 8.20 m2 on the ground. This dataset contains 1000 samples and we use 800 samples for training and 100 for validation and 100 for test.
V-B2 Semantic Segmentation Drone dataset
The Semantic Drone Segmentation dataset contains 400 high-resolution nadir-view images of urban residential scenes, captured at altitudes between 5 m and 30 m above ground and annotated with detailed pixel-level semantic labels covering 24 object categories. In our experiments, we use 340 images for training, 32 for validation, and 28 for testing.
V-C Evaluation with Different Convolution Approaches
V-C1 Single orientation rotation convolution
Given that single-orientation rotation convolution collapses to standard convolution, we first conduct a direct comparison between our scatter-based convolution API and cuDNN’s standard convolution API. The corresponding results are reported in Appendix A.
We compare our approach with cuDNN as well as the state-of-the-art ORN [orn] and E2CNN [e2cnn] frameworks, both of which provide APIs for achieving rotation invariance. Tables I and II report the performance of cuDNN, ORN, E2CNN, and the proposed scatter-based convolution under the single-orientation setting. The observed accuracy variation is within $\pm 0.1\%$. Since single-orientation rotation convolution degenerates to standard convolution, all methods achieve nearly identical segmentation accuracy. Nevertheless, notable differences arise in computational efficiency. On the plant dataset with 256×256 inputs, E2CNN achieves the lowest training time and energy consumption, while scatter-based convolution also demonstrates competitive efficiency, training approximately 9.2% faster than cuDNN and reducing energy consumption by about 6.9%. At 1024×1024 resolution, scatter-based convolution achieves the best efficiency, reducing training time by roughly 7.0% and energy usage by 7.6% compared with cuDNN. Similar trends are observed on the drone dataset: at 256 resolution, E2CNN attains the highest accuracy and lowest computational cost, whereas scatter-based convolution remains consistently more efficient than cuDNN. At 1024 resolution, scatter-based convolution outperforms all baselines in both training time (7.1% reduction) and energy consumption (7.9% reduction). These results indicate that while E2CNN is particularly efficient for smaller input resolutions, the proposed scatter-based convolution provides consistent and scalable efficiency gains—especially at higher resolutions—without sacrificing segmentation accuracy, as it avoids explicit data lowering and reduces memory traffic through direct scatter accumulation. Representative segmentation results are visulized in Fig. 10 and Fig. 11.
TABLE I: Plant Data Segmentation Using Various Single-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 8604.0 0.379 73.59
ORN conv 7758.0 0.362 73.63
E2CNN conv 7116.0 0.315 73.73
scatter based conv 7808.0 0.353 73.62
1024 cudnn based conv 26153.0 1.758 71.32
ORN conv 25538.0 1.719 71.38
E2CNN conv 27405.0 2.213 71.66
scatter based conv 24317.0 1.625 71.28
Figure 10: Visualization of segmentation results on plant data.
TABLE II: Semantic drone segmentation Using Various Single-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 3242.2 0.1408 90.62
ORN conv 2945.1 0.1359 90.67
E2CNN conv 2705.7 0.1186 90.73
scatter based conv 2925.6 0.1325 90.65
1024 cudnn based conv 9865.5 0.6618 88.71
ORN conv 9726.3 0.6498 88.74
E2CNN conv 10392.7 0.8372 88.78
scatter based conv 9166.9 0.6094 88.69
Figure 11: Visualization of segmentation results on Semantic Segmentation Drone dataset.
V-C2 Symetric four orientation rotation convolution
Tables III and IV summarize the results for the symmetric four-orientation configuration. Relative to single-orientation convolution, accuracy increases by about 2%, with E2CNN achieving the highest accuracy, reflecting the benefit of four rotation modeling. The computational advantages of scatter-based convolution become more pronounced under this setting. On the plant dataset, scatter reduces training time by approximately 45.9% at 256 resolution and 28.8% at 1024 resolution, while lowering energy consumption by about 38.4% and 18.6%, respectively. On the drone dataset, scatter provides comparable benefits, achieving about 46.7% faster training at 256 resolution and 32.9% at 1024 resolution, with corresponding energy reductions of roughly 39.1% and 22.8%. These results indicate that symmetric four-orientation convolutions achieve strong segmentation accuracy with manageable computational cost. The scatter based convolution provides the most favorable accuracy–efficiency profile, benefiting from reuse of intermediate computations across symmetric rotations. This enables substantial reductions in training time and energy consumption while maintaining segmentation accuracy.
TABLE III: Plant Data Segmentation Using Various Four-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 20234.6 0.943 75.64
ORN conv 20154.2 0.921 75.69
E2CNN conv 18575.8 0.783 75.71
scatter based conv 10962.6 0.581 75.61
1024 cudnn based conv 82081.0 5.564 72.92
ORN conv 81706.5 5.623 72.98
E2CNN conv 76175.0 5.118 73.05
scatter based conv 58433.9 4.532 72.95
TABLE IV: Semantic drone segmentation Using Various Four-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 7713.1 0.3570 91.88
ORN conv 7651.6 0.3501 91.93
E2CNN conv 6908.2 0.2943 92.06
scatter based conv 4109.6 0.2175 91.88
1024 cudnn based conv 30936.5 2.103 90.39
ORN conv 31025.2 2.203 90.41
E2CNN conv 28925.6 1.946 90.62
scatter based conv 20761.8 1.623 90.43
V-C3 Symetric eight orientation rotation convolution
Tables V and VI report the eight-orientation results. Increasing the number of modeled orientations leads to further accuracy improvements, especially for the plant dataset. Because convolutional pooling over eight orientations produces twice the feature-map dimensionality of the four-orientation configuration. However, it imposes substantially greater memory and computational demands on conventional implementations. The scatter based convolution continues to provide substantial efficiency benefits. On the plant dataset, scatter is approximately 45.4% faster than cuDNN at 256 resolution and about 34.5% faster at 1024 resolution, while reducing energy usage by approximately 41.2% and 21.7%, respectively. On the drone dataset, scatter achieves around 56.9% faster training at 256 resolution and 32.7% faster at 1024 resolution, with energy savings of about 44.3% and 22.7%, respectively.
The proposed scatter-based approach differs from E2CNN in how rotation is modeled. While E2CNN learns orientation-dependent expansion coefficients, the scatter-based method constructs steerable filters from a set of learnable base filters, as described in Section III-D, where orientation dependence is introduced analytically. By enforcing orthogonality and equal-magnitude constraints on these base filters, the formulation approximates rotation invariance, leading to accuracy trends distinct from E2CNN.
The experimental results demonstrate that as rotation resolution increases, rotation invariant convolutions become more accurate, and the scatter based implementation remains more efficient than ORN and E2CNN. ORN shows smaller accuracy improvements in the eight-orientation setting because it does not provide an API for orientation-subgroup pooling; in our implementation, its group pooling operation aggregates all eight orientations into a single pooled feature. As accuracy continues to improve, conventional rotation-invariant convolutions incur rapidly increasing memory and computation overhead due to expanded feature-map dimensionality. This effect is particularly evident for E2CNN, where group representations lead to larger intermediate tensors. By contrast, the proposed scatter-based implementation mitigates this issue by consolidating rotation handling within a single convolution pass, thereby maintaining scalability as orientation resolution increases.
TABLE V: Plant Data Segmentation Using Various Eight-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 36866.0 1.675 78.69
ORN conv 36188.0 1.612 76.13
E2CNN conv 33408.0 1.561 78.15
scatter based conv 20133.0 0.9844 78.64
1024 cudnn based conv 100269.79 6.8162 75.86
ORN conv 95240.30 6.5967 74.19
E2CNN conv 84227.06 5.8908 75.57
scatter based conv 65703.51 5.3363 75.82
TABLE VI: Semantic drone segmentation Using Various Eight-Orientation Rotation-Invariant Convolutions
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 cudnn based conv 14002.8 0.6345 93.98
ORN conv 13798.3 0.6132 92.76
E2CNN conv 12682.3 0.5894 93.67
scatter based conv 6042.28 0.3533 93.91
1024 cudnn based conv 41836.42 2.8440 92.44
ORN conv 41382.05 2.8402 91.20
E2CNN conv 36342.18 2.4986 91.99
scatter based conv 28137.53 2.1996 92.40
V-C4 Symetric sixteen orientation rotation convolution
In preliminary experiments, we found that modeling sixteen orientations in UNet leads to very large intermediate feature tensors and impractical memory and compute overhead. Compared to implementations that process each rotation separately, our scatter-based approach performs symmetric four-rotation convolution and pooling in a single pass within GPU on-chip memory, reducing the number of intermediate feature maps by a factor of four. As a result, our scatter-based implementation is the only approach capable of supporting a full sixteen-orientation convolution and pooling within practical memory and runtime constraint. As shown in Tables VII and VIII, the scatter-based approach exhibits improved accuracy when employing sixteen orientations relative to the eight-orientation configuration. Because convolutional pooling over sixteen orientations generates twice the feature-map dimensionality of the eight-orientation configuration.
TABLE VII: Plant Data Segmentation Using Sixtee-Orientation Rotation-Invariant Convolution
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 scatter based conv 31208.3 1.555 79.27
1024 scatter based conv 103612.2 6.685 76.33
TABLE VIII: Semantic drone segmentation Using Sixteen-Orientation Rotation-Invariant Convolution
Image size Convolution applied Training time (s) Energy (kWh) Test accuracy (%)
256 scatter based conv 9982.3 0.513 94.39
1024 scatter based conv 41905.6 3.101 92.87
V-D Post-Segmentation Processing and Visualization
The rotation-invariant U-Net segmentation network was trained to segment plant species in the imagery, and the plant-species segmentation masks from all frames (see Fig. 12) were subsequently georeferenced and mosaicked into a seamless orthomosaic. The resulting species‑labeled mosaic was exported to Google Earth Pro for visualization (Fig. 13) and for analysis of species distribution across the field. As shown in Fig. 14, the plant area consists of 51.41% clover leaf, 37.89% grass, 3.13% dock leaf, and 7.56% weeds and other vegetation, indicating clover as the dominant ground cover followed by grass, with minimal presence of dock leaves and miscellaneous weeds.
Figure 12: Original P1 RGB imagery (left) and corresponding segmentation masks (right).
Figure 13: Geo-referenced P1 RGB imagery mosaic (Top) and corresponding Geo-referenced mask mosaic (Bottom) displayed in Google Earth Pro.
(a) Visualization of DSM (white) and Soil segmentation mask (black inside white DSM).
(b) Visualization of Clover leaf segmentation mask.
(c) Visualization of Grass segmentation mask.
(d) Visualization of Dock leaf segmentation mask
(e) Visualization of weeds and other plant segmentation mask.
Figure 14: Visualization of individual species masks
VI Conclusion
In conclusion, replacing cuDNN’s dense kernels with our single-pass scatter convolution accelerates both standard (single-orientation) and rotation-invariant layers without affecting accuracy. For standard convolutions, the scatter kernel lowers memory pressure and computational load, reducing training time and energy consumption on small- and medium-sized feature maps with deep channels. When rotations are required it fuses all filters into one pass, yielding even larger speed-and-energy gains. Results on the Semantic Drone and plant-segmentation benchmarks confirm consistent improvements across input sizes, establishing scatter convolution as a practical drop-in upgrade for modern CNNs.
Acknowledgments
This work is funded by the UK Natural Environment Research Council under project FLORA-SAGE - Federated Learning Optimised for Remote Assessment of Species in Agricultural Grassland Ecosystems (grant award number UKRI053) and UKRI Digital Dairy Value-Chain for South-West Scotland and Cumbria (Strength in Places Fund) (grant award number 99890)
Appendix A Performance of Scatter-Based Convolution Compared to cuDNN Convolutions
In the single-rotation case, the rotation-invariant convolution reduces to a standard convolution. We evaluate our scatter-based convolution (single orientation) through a series of experiments and benchmark it against NVIDIA’s cuDNN convolution library, which is widely used in deep-learning applications [cuDNNsdk]. In our implementation, the fastest cuDNN kernel is automatically selected by the cuDNN autotuner. The GPU used for evaluation is NVIDIA GeForce RTX 3080 Ti with 16GB video memory. The version of CUDA 12.3 and cuDNN 8.9.6.50 are installed in a Windows 10 system.
48163264128256
256
48163264128512
256
481632641281024
256
00.10.10.20.20.30.3 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (a) $4\times 4$ input
48163264128256
256
48163264128512
256
481632641281024
256
00.20.20.40.40.60.60.80.8 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (b) $8\times 8$ input
48163264128256
256
48163264128512
256
481632641281024
256
01122 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (c) $16\times 16$ input
48163264128256
256
48163264128512
256
481632641281024
256
022446688 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (d) $32\times 32$ input
48163264128256
256
48163264128512
256
481632641281024
256
01010202030304040 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (e) $64\times 64$ input
48163264
64
481632128
64
481632256
64
02244 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (f) $128\times 128$ input
481632
64 64
481632
128 64
481632
256 64
055101015152020 Primary X-axis: Number of filters Secondary X-axis: Number of input channels Run time (ms)CUDNN Fastest ConvConv With Scatter Operation
Fig.15 (g) $256\times 256$ input
Figure 15: Performance comparison of cuDNN (blue) and the proposed scatter convolution (orange) across seven input resolutions.
NVIDIA’s cuDNN offers eight distinct forward-convolution implementations, and for every experiment we measure the fastest of those eight via cuDNN’s autotuner. As shown in Fig.15 (a)-(g), across the seven plots (input sizes $4\to 256$) our scatter-based kernel still dominates that “best-of-eight” cuDNN baseline in the majority of layer configurations. Counting every (resolution $\times$ input-channel $\times$ filter) triplet, the scatter variant is quicker in most cases, with typical speed-ups of $1.1\times$$1.6\times$ and a peak of $\approx 2\times$ for the very small $4\times 4$ activations. The gain is most pronounced whenever the spatial footprint is modest ($\leq 32\times 32$) or the channel–filter product is under $\sim 16\text{\,K}$, where kernel-launch latency and global-memory traffic dominate the cost of convolution; the single-pass scatter design removes the data-lowering staging step and sustains higher memory-bandwidth utilisation in that regime. As the resolution grows beyond $64\times 64$ and the workload becomes compute-bound, cuDNN paths occasionally pull ahead, yet even at $256\times 256$ our kernel still wins for 4 of the 9 channel–filter pairs tested. In practice, this means the scatter kernel is the better choice for mid‑to‑deep layers where spatial dimensions have shrunk and channel depth has increased, while cuDNN remains preferable only for the first one or two encoder layers of very high-resolution networks.
[Uncaptioned image] Manduhu Manduhu received the M.Sc. and Ph.D. degrees in information engineering from Hiroshima University, Japan, in 2010 and 2013, respectively. He is currently a PostDoc researcher at the Drone Systems Lab of the University of the West of Scotland. His research interests include parallel algorithms, physics‑based simulation, and deep learning for object detection and segmentation.
[Uncaptioned image] Alexander Dow received his B.S. in sound engineering and production from Birmingham City University, Birmingham, UK in 2016 and his M.S. in digital signal and image processing from the University of Sussex, Brighton, UK in 2020. He is currently pursuing his Ph.D. in drone-based computer vision at the University of the West of Scotland, Lanarkshire, Scotland. He currently works as a Post-Doctoral Research Assistant in the Drone Systems Lab at the University of the West of Scotland. His research interests include computer vision, artificial intelligence on embedded systems, LiDAR and hyperspectral imaging.
[Uncaptioned image] Gerard Dooly received the B.Eng. degree from the Electronic and Computer Engineering Department, University of Limerick, in 2003, and the Ph.D. degree from the Optical Fibre Sensors Research Centre, University of Limerick, in 2008, on the topic “An Optical Fibre Sensor for the Measurement of Hazardous Emissions from Land Transport Vehicles.” For over ten years, he has worked extensively in field robotics at UL. His research interests include real-time 3-D reconstruction, machine vision, machine learning, optical fibre sensors, subsea structural health monitoring, teleoperation, and automated docking and intervention. He is involved in robotics for harsh environments in offshore setting and is developing systems to address beyond visual line of sight operations for UAS. He is focused on the design and development of robotics and has engaged in numerous field operations and survey missions both in Ireland and on the continent. Some of his recent offshore operations involved environmental sensing, anti-mine countermeasure ops, remote UAS for incident response, archaeological survey, and hybrid long range UAS technologies.
[Uncaptioned image] James Riordan received the B.Eng. degree in electronic engineering specializing in aircraft simulation systems and the Ph.D. degree in real-time processing of acoustic signals from the University of Limerick, Ireland. Currently, he is a Full Professor with the University of the West of Scotland, U.K., where he is also the Director of the Drone Systems Laboratory. He is also a Principal Investigator of multiple research projects funded by European Commission and U.K. Research and Innovation. His research interests include artificial intelligence, computer vision, and sensing methods to extend the safe and sustainable application of autonomous vehicles in land, air, and sea.